OpenStudy (anonymous):

Dan poured 5 cups of water into a rectangular container measuring 8 inches by 6 inches by 7 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. (3 points) Part B: How many cubic inches of water was poured into the container? (3 points) Part C: What was the height of the water in the container? Show your work. (4 points) Measurement is given as length by width by height. please help

OpenStudy (anonymous):

A. volume of container=8*6*7/2.54=132.28 cubic inch. B.Required volume =5*14.44=72.2 cubic inch C. Let h be the height 8*6*h=72.2 h=72.2/48=1.5 inch.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw